A daughters present age is half the present age of her father. 10 years ago the father was thrice as old as his daughter. What are their present ages?
Answer by Alan3354(69443)

A daughters present age is half the present age of her father. 10 years ago the father was thrice as old as his daughter. What are their present ages?
-------------
F = 2D
F - 10 = 3(D - 10)
------
Sub for F
2D - 10 = 3D - 30
D = 20